# MS-ISSM: Objective Quality Assessment of Point Clouds Using Multi-scale Implicit Structural Similarity

### v1

@@ IV-D ResGrouped-MLP Regression
We propose the ResGrouped-MLP, a hierarchical deep learning framework designed to robustly map hand-crafted point cloud features to subjective quality scores (MOS). Addressing the limitations of flat networks, we adopt a ”Split-Transform-Merge” architecture to preserve the distinct physical semantics of features. The framework integrates a novel Log-Modulus preprocessing and a multi-scale attention mechanism, as shown in Fig. 4.
Figure 5: Data Pipeline and Validation Strategy
Log-Modulus Preprocessing: Statistical features extracted from point clouds often exhibit a heavy-tailed distribution, where standard Z-score normalization can lead to gradient instability and poor convergence. To address this, we introduce the Log-Modulus transformation before normalization. Unlike simple logarithmic transforms, this method handles both positive and negative values while compressing the dynamic range:
$\tilde{\rm{x}}=\text{sign}(\rm{x})\cdot\ln(1+|\rm{x}|)\text{,}$ (16)
@@
The ResGrouped-MLP Architecture: Point cloud quality perception relies on the interplay between feature scales (High, Medium, Low) and attribute channels (Luma, Chroma, Geometry). Simply concatenating these features risks losing their distinct physical meanings. Therefore, we design a hierarchical architecture composed of three key stages:
a) Deep Grouped Encoders: Instead of a generic fully connected layer, we treat each scale-channel pair as an independent group. We employ Residual Blocks for feature encoding, formulated as $x+\mathcal{F}(x)$, where $\mathcal{F}$ represents a dual-layer perceptron with Batch Normalization and SiLU activation. This grouped design isolates the specific distortion characteristics of each channel, preventing information interference at early stages while mitigating the vanishing gradient problem.
@@
$\mathbf{F}{scale}^{\prime}=\mathbf{F}{scale}\otimes\sigma(\text{MLP}(\mathbf{F}{scale}))\text{,}$ (17)
@@
where $\lambda_{1}$ and $\lambda_{2}$ are weighting parameters used to balance the optimization objectives.
Implementation and Validation: The model is trained using the AdamW optimizer with a weight decay of $10^{-2}$ and a Cosine Annealing scheduler for 80 epochs (batch size 32). For validation, we employ a rigorous Repeated Random Shuffle Split strategy (5 rounds), partitioning the dataset into 60% training and 40% testing sets. This large test ratio ensures the model is evaluated on a substantial amount of unseen data, demonstrating its generalization capability.
@@ V Experimental Evaluations
TABLE I: Performance comparison of different PCQA metrics. The best and second-best are highlighted in red and blue, respectively.
@@
Method Crit. Datasets M-PCCD [43] WPC [40] ICIP [44] SJTU [10] ALL PSNR-p2p [7] PLCC 0.408 0.318 0.566 0.592 0.404 SROCC 0.475 0.297 0.579 0.662 0.398 KROCC 0.354 0.202 0.435 0.493 0.278 RMSE 0.446 0.448 0.331 0.376 0.504 PSNR-p2pl [7] PLCC 0.422 0.250 0.560 0.534 0.391 SROCC 0.489 0.229 0.583 0.602 0.380 KROCC 0.368 0.159 0.458 0.444 0.269 RMSE 0.446 0.457 0.342 0.400 0.504 PSNR-YUV [7] PLCC 0.603 0.274 0.769 0.535 0.439 SROCC 0.620 0.260 0.777 0.553 0.414 KROCC 0.442 0.177 0.580 0.385 0.280 RMSE 0.313 0.353 0.212 0.363 0.396 GraphSIM [25] PLCC 0.912 0.597 0.889 0.854 0.716 SROCC 0.936 0.556 0.888 0.850 0.704 KROCC 0.786 0.412 0.723 0.649 0.524 RMSE 0.246 0.223 0.193 0.194 0.205 MS-GraphSIM [26] PLCC 0.893 0.614 0.892 0.888 0.737 SROCC 0.922 0.581 0.890 0.882 0.726 KROCC 0.761 0.453 0.727 0.692 0.560 RMSE 0.250 0.224 0.197 0.200 0.208 PCQM [14] PLCC 0.913 0.310 0.869 0.576 0.467 SROCC 0.916 0.378 0.956 0.772 0.606 KROCC 0.748 0.306 0.830 0.589 0.470 RMSE 0.288 0.241 0.229 0.243 0.242 PSSIM [27] PLCC 0.592 0.497 0.675 0.759 0.595 SROCC 0.716 0.424 0.778 0.721 0.579 KROCC 0.538 0.300 0.601 0.525 0.414 RMSE 0.262 0.227 0.200 0.207 0.216 MS-PSSIM [31] PLCC 0.814 0.658 0.839 0.876 0.703 SROCC 0.927 0.655 0.841 0.900 0.705 KROCC 0.774 0.478 0.669 0.721 0.508 RMSE 0.264 0.224 0.212 0.208 0.221 TCDM [29] PLCC 0.921 0.795 0.958 0.924 0.847 SROCC 0.947 0.807 0.964 0.924 0.861 KROCC 0.808 0.610 0.863 0.760 0.664 RMSE 0.249 0.211 0.262 0.190 0.204 FRSVR [36] PLCC 0.754 0.243 0.772 0.549 0.438 SROCC 0.818 0.274 0.761 0.679 0.480 KROCC 0.621 0.175 0.587 0.478 0.317 RMSE 0.206 0.276 0.160 0.198 0.221 PHM [30] PLCC 0.935 0.813 0.831 0.892 0.796 SROCC 0.958 0.849 0.953 0.896 0.852 KROCC 0.829 0.660 0.833 0.717 0.656 RMSE 0.275 0.238 0.211 0.216 0.227 MS-ISSM (Ours) PLCC 0.944 0.931 0.925 0.886 0.913 SROCC 0.949 0.935 0.953 0.863 0.914 KROCC 0.811 0.774 0.833 0.701 0.751 RMSE 0.224 0.203 0.193 0.214 0.188
@@
TABLE II: Performance comparison of classic and SOTA metrics on different distortion types. The best and second-best are highlighted in red and blue, respectively.
Distortion types
Metrics Criterias Ds Ns Oc Mx Tc Vc Rank
PLCC 0.275 0.589 0.412 0.633 0.426 0.207 10.17
SROCC 0.339 0.609 0.324 0.646 0.437 0.293 10.33
KROCC 0.233 0.442 0.229 0.468 0.307 0.199 10.50
PSNR-p2p RMSE 0.411 0.308 0.322 0.288 0.345 0.326 11.17
PLCC 0.208 0.607 0.421 0.635 0.440 0.240 9.33
SROCC 0.180 0.622 0.326 0.598 0.435 0.334 10.33
KROCC 0.122 0.453 0.230 0.425 0.311 0.230 10.17
PSNR-p2pl RMSE 0.440 0.306 0.318 0.296 0.324 0.323 10.83
PLCC 0.655 0.476 0.567 0.604 0.276 0.452 10.00
SROCC 0.646 0.507 0.613 0.680 0.271 0.448 10.00
KROCC 0.443 0.368 0.419 0.522 0.192 0.310 10.17
PSNR-YUV RMSE 0.287 0.310 0.317 0.299 0.381 0.230 10.83
PLCC 0.939 0.888 0.856 0.859 0.494 0.630 4.50
SROCC 0.891 0.894 0.862 0.850 0.442 0.603 6.00
KROCC 0.675 0.710 0.678 0.640 0.328 0.425 5.83
GraphSIM RMSE 0.247 0.173 0.214 0.204 0.243 0.167 3.00
PLCC 0.935 0.907 0.882 0.903 0.507 0.652 3.33
SROCC 0.903 0.910 0.882 0.869 0.455 0.640 4.33
KROCC 0.712 0.733 0.699 0.667 0.337 0.452 4.50
MS-GraphSIM RMSE 0.254 0.177 0.217 0.209 0.243 0.169 4.33
PLCC 0.120 0.877 0.607 0.429 0.293 0.594 9.17
SROCC 0.159 0.916 0.896 0.852 0.387 0.568 7.00
KROCC 0.036 0.743 0.707 0.667 0.326 0.407 6.33
PCQM RMSE 0.291 0.218 0.255 0.249 0.262 0.182 9.17
PLCC 0.697 0.629 0.633 0.738 0.218 0.234 9.00
SROCC 0.719 0.694 0.729 0.894 0.290 0.358 8.17
KROCC 0.563 0.514 0.524 0.719 0.215 0.246 8.17
PSSIM RMSE 0.286 0.212 0.246 0.240 0.260 0.179 8.00
PLCC 0.844 0.866 0.798 0.893 0.654 0.551 5.83
SROCC 0.874 0.879 0.839 0.911 0.627 0.544 5.33
KROCC 0.672 0.683 0.647 0.737 0.444 0.378 5.50
MS-PSSIM RMSE 0.278 0.185 0.234 0.220 0.235 0.169 5.67
PLCC 0.898 0.909 0.899 0.932 0.834 0.647 2.67
SROCC 0.873 0.906 0.901 0.935 0.845 0.657 3.33
KROCC 0.671 0.732 0.723 0.779 0.654 0.477 3.17
TCDM RMSE 0.261 0.170 0.219 0.205 0.222 0.165 3.00
PLCC 0.809 0.522 0.733 0.560 0.408 0.417 9.00
SROCC 0.644 0.611 0.752 0.594 0.449 0.453 9.00
KROCC 0.465 0.432 0.552 0.404 0.292 0.319 9.50
FRSVR RMSE 0.265 0.178 0.180 0.211 0.235 0.164 3.50
PLCC 0.906 0.913 0.824 0.896 0.763 0.663 3.17
SROCC 0.897 0.916 0.910 0.909 0.790 0.664 2.33
KROCC 0.693 0.749 0.735 0.742 0.599 0.468 2.50
PHM RMSE 0.277 0.195 0.236 0.229 0.246 0.173 6.83
PLCC 0.900 0.919 0.892 0.920 0.930 0.914 1.83
SROCC 0.934 0.924 0.905 0.886 0.922 0.913 1.83
KROCC 0.737 0.756 0.737 0.704 0.757 0.754 1.67
MS-ISSM (Ours) RMSE 0.228 0.166 0.214 0.208 0.216 0.160 1.67
V-A Datasets and PCQA Metrics under Comparison
To verify the performance of the proposed method across different types of distortions, we use four point cloud subjective datasets for assessment. These datasets include: SJTU [10], WPC [40], M-PCCD[43], and ICIP [44]. We also combine these four datasets into a comprehensive dataset, ALL, to validate the stability and reliability of objective metrics. It is important to note that the subjective rating scales differ across these datasets. To address this, we map the subjective scores from all four datasets onto a common scale of [0, 1] for consistency. In addition, this paper compares the proposed algorithm with 11 classic and SOTA PCQA metrics.
V-B Evaluation Criteria
@@ V-C Performance Comparison
We evaluate the performance of various PCQA metrics using different datasets. The overall evaluation results for each PCQA metric across these datasets are presented in Table I. To facilitate direct comparison, the best and second-best performing metrics for each dataset are highlighted in red and blue, respectively. Specifically, the proposed method ranks highest on the WPC dataset compared to the other 11 PCQA metrics. On the M-PCCD dataset, the proposed method achieved first place in PLCC, and second place in SROCC, KROCC, and RMSE. On the ICIP dataset, it performs second only to TCDM. Although the proposed method does not achieve the best performance on the SJTU dataset, it closely trails the top-performing methods. On the SJTU dataset, MS-ISSM yields results of [0.886, 0.863, 0.701, 0.214], compared with the best-performing method, TCDM, which achieves [0.924, 0.924, 0.760, 0.190]. Notably, on the combined ALL dataset comprising all four datasets, the proposed method demonstrates superior performance with [0.913, 0.914, 0.751, 0.188], closely followed by TCDM with [0.847, 0.861, 0.664, 0.204]. Overall, the proposed method demonstrates superior performance.
@@
Table II presents the performance of the proposed method and other metrics across various distortion types, including octree-based compression distortion (Oc), video-based compression distortion (Vc), trisoup-based compression distortion (Tc), noise distortion (Ns), downsampling distortion (Ds), and mixing distortion (Mix). As with previous tables, the best and second-best-performing metrics for each distortion type are highlighted in red and blue, respectively. In general, the proposed MS-ISSM exhibits dominating performance in Downsampling, Noise, Trisoup-based compression, and Video-based compression distortions. Specifically, it achieves the top rank across all four indicators (PLCC, SROCC, KROCC, and RMSE) for these categories. For Octree-based compression, MS-ISSM secures the best KROCC score, with other metrics closely following the top performer.
@@
Regarding Mixing distortion, MS-ISSM performs slightly lower than TCDM. This is primarily because the mixing distortion samples originate entirely from the SJTU dataset, on which TCDM’s parameters were explicitly fitted, granting it a distributional advantage in this specific scenario. However, in the more challenging Video-based compression scenario, MS-ISSM achieves a significant lead with performance scores of [0.914, 0.913, 0.754, 0.160], far surpassing the second-best method, which only reaches [0.663, 0.664, 0.477, 0.164].
@@
When ranking all methods across different distortion types, the proposed method achieves the best overall performance in terms of PLCC, SROCC, KROCC, and RMSE. This generalization capability confirms the effectiveness of our design: (1) Robustness of Implicit Surface Reconstruction: The RBF-based implicit representation reconstructs continuous surfaces to resolve sparsity and geometric jitter. This bypasses discrete point-to-point matching errors, yielding superior stability in V-PCC and compression distortions. (2) Comprehensiveness of Multi-scale Strategy: Our hierarchical approach captures both global structural shifts (Low Scale) from downsampling and local high-frequency artifacts (High Scale) from compression. (3) Adaptive Non-linear Mapping: The ResGrouped-MLP combines Log-Modulus transformation for distribution rectification with Channel-wise Attention for adaptive feature weighting, ensuring robust prediction in mixed distortion scenarios.
@@
Additionally, to further compare the performance of the MS-ISSM method and the RBFIM method in terms of compression distortion, we evaluated datasets containing compression distortion from the three aforementioned datasets. Based on the type of compression distortion, the datasets were divided into G-PCC compression distortion and V-PCC compression distortion. The experimental results in Table III reveal that RBFIM performs better on the ICIP G-PCC compression distortion portion. However, MS-ISSM significantly outperforms RBFIM in the V-PCC portion of both M-PCCD and WPC, and performs better across all compression distortion types. This indicates that the multi-scale implicit feature method exhibits stronger generalization and robustness.
@@
TABLE III: Comparison of the RBFIM and MS-ISSM methods under different compression distortions.
@@
ICIP-GPCC PLCC 0.993 0.835
SROCC 0.971 0.841
KROCC 0.870 0.610
RMSE 0.020 0.194
ICIP-VPCC PLCC 0.969 0.965
SROCC 0.976 0.996
KROCC 0.895 0.995

### v2

@@ IV-D ResGrouped-MLP Regression
We propose the ResGrouped-MLP, a hierarchical deep learning framework designed to robustly map hand-crafted point cloud features to subjective quality scores. Addressing the limitations of flat networks, we adopt a ”Split-Transform-Merge” architecture to preserve the distinct physical semantics of features. The framework integrates a novel Log-Modulus preprocessing and a multi-scale attention mechanism, as shown in Fig. 4.
Figure 5: Preprocessing pipeline of hand-crafted features
Log-Modulus Preprocessing: Statistical features extracted from point clouds often exhibit a heavy-tailed distribution, where standard Z-score normalization can lead to gradient instability and poor convergence. To address this, we selected the Log-Modulus transformation to handle signed data and compress heavy-tailed distributions caused by local distortions. This parameter-free operation improves network training stability as follows:
${\rm{x}^{\prime}}=\text{sign}(\rm{x})\cdot\ln(1+|\rm{x}|)\text{,}$ (17)
@@
The ResGrouped-MLP Architecture: Point cloud quality perception relies on the interplay between feature scales (High, Medium, Low) and attribute channels (luma, chroma, curvature). Simply concatenating these features risks losing their distinct physical meanings. Therefore, we design a hierarchical architecture composed of three key stages:
a) Deep Grouped Encoders: Instead of a generic fully connected layer, we treat each scale-channel pair as an independent group. We employ Residual Blocks for feature encoding, formulated as $\rm{x}+\mathcal{F}(\rm{x})$, where $\mathcal{F}()$ represents a dual-layer perceptron with Batch Normalization and SiLU activation. This grouped design isolates the specific distortion characteristics of each channel, preventing information interference at early stages while mitigating the vanishing gradient problem.
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{}\mathbf{f}{scale}^{\prime}=\mathbf{f}{scale}\otimes\sigma(\text{MLP}(\mathbf{f}{scale}))$$\text{,}$ (18)
@@
where $\lambda_{1}$ and $\lambda_{2}$ are weighting parameters used to balance the optimization objectives. And, these values are set to 0.2 and 0.2, respectively. Such a configuration ensures that the PLCC loss provides strong linear correlation guidance and the Margin Ranking loss penalizes monotonicity violations without overwhelming the primary MSE regression.
Implementation and Validation: The model is trained using the AdamW optimizer with a weight decay of $10^{-2}$ and a Cosine Annealing scheduler for 80 epochs (batch size 32). The detailed hyperparameter configurations, including the grouped encoders, scale-wise attention and global regressor, are summarized in Table LABEL:tab:config.
TABLE I: Detailed Configurations of the ResGrouped-MLP.
Module Sequential Operations
grouped encoders FC (64), BN, SiLU, ResBlock (dropout = 0.1)
scale-wise attention FC (48), SiLU, FC (192), sigmoid
global regressor FC (128), BN, SiLU, ResBlock (dropout = 0.2), FC (64), SiLU, FC (1), sigmoid
@@ V Experimental Evaluations
TABLE II: Performance comparison of different PCQA metrics. The best, second-best, and third-best among the primary methods are highlighted in red, blue, and underlined, respectively.
@@
Method Crit. Datasets Rank SJTU [10] ICIP [45] M-PCCD [44] WPC [40] PSNR-p2p [7] PLCC 0.460 0.862 0.395 0.260 12.25 SROCC 0.398 0.853 0.239 0.008 12.00 KROCC 0.298 0.590 0.196 0.002 12.75 RMSE 2.130 0.659 1.250 22.137 12.50 PSNR-p2pl [7] PLCC 0.433 0.898 0.395 0.260 11.75 SROCC 0.415 0.843 0.232 0.053 12.00 KROCC 0.303 0.606 0.201 0.034 11.75 RMSE 2.162 0.642 1.250 22.137 11.75 PSNR-YUV [7] PLCC 0.550 0.779 0.603 0.387 11.50 SROCC 0.555 0.777 0.620 0.384 11.50 KROCC 0.385 0.580 0.442 0.263 11.50 RMSE 2.003 0.712 1.085 21.139 11.50 PSSIM [27] PLCC 0.769 0.886 0.712 0.500 10.00 SROCC 0.717 0.877 0.716 0.424 9.75 KROCC 0.522 0.701 0.538 0.300 9.75 RMSE 1.533 0.653 0.956 19.854 10.25 PCQM [14] PLCC 0.807 0.948 0.903 0.563 6.75 SROCC 0.774 0.956 0.916 0.545 6.50 KROCC 0.590 0.830 0.748 0.432 7.25 RMSE 1.418 0.363 0.584 18.947 7.00 MS-PSSIM [31] PLCC 0.844 0.929 0.768 0.551 7.50 SROCC 0.842 0.906 0.777 0.541 7.50 KROCC 0.645 0.771 0.602 0.381 7.25 RMSE 1.286 0.612 0.872 19.132 8.25 GraphSIM [25] PLCC 0.860 0.901 0.921 0.690 6.00 SROCC 0.843 0.888 0.936 0.679 6.25 KROCC 0.641 0.723 0.786 0.493 6.50 RMSE 1.223 0.492 0.529 16.597 6.00 MS-GraphSIM [26] PLCC 0.897 0.902 0.906 0.702 5.75 SROCC 0.874 0.890 0.922 0.701 5.75 KROCC 0.682 0.727 0.761 0.517 6.00 RMSE 1.058 0.491 0.575 16.324 5.50 PHM [30] PLCC 0.901 0.867 0.945 0.838 4.50 SROCC 0.880 0.831 0.948 0.832 4.75 KROCC 0.695 0.701 0.807 0.639 4.50 RMSE 1.038 0.566 0.445 12.509 4.00 TCDM [29] PLCC 0.931 0.942 0.937 0.805 3.50 SROCC 0.912 0.935 0.945 0.803 3.75 KROCC 0.740 0.803 0.801 0.604 3.75 RMSE 0.875 0.381 0.475 13.607 4.00 FRSVR (K-fold splicing) [36] PLCC 0.822 0.957 0.886 0.825 5.00 SROCC 0.806 0.954 0.904 0.827 5.25 KROCC 0.601 0.843 0.783 0.620 4.50 RMSE 1.333 0.341 0.341 12.786 3.25 PointPCA (K-fold splicing) [34] PLCC 0.819 0.942 0.912 0.806 5.25 SROCC 0.786 0.946 0.946 0.807 4.75 KROCC 0.592 0.837 0.827 0.607 4.25 RMSE 1.341 0.375 0.375 13.456 4.75 MS-ISSM (K-fold splicing) PLCC 0.906 0.964 0.958 0.855 1.25 SROCC 0.897 0.967 0.961 0.846 1.25 KROCC 0.698 0.852 0.853 0.666 1.25 RMSE 1.224 0.283 0.351 11.736 2.25 FRSVR (K-fold average) [36] PLCC 0.858 0.976 0.916 0.842 - SROCC 0.823 0.969 0.929 0.836 - KROCC 0.622 0.893 0.778 0.647 - RMSE 1.279 0.279 0.432 12.143 - PointPCA (K-fold average) [34] PLCC 0.847 0.976 0.956 0.812 - SROCC 0.842 0.974 0.964 0.815 - KROCC 0.625 0.855 0.855 0.614 - RMSE 1.293 0.269 0.339 13.078 - MS-ISSM (K-fold average) PLCC 0.925 0.968 0.966 0.867 - SROCC 0.911 0.976 0.971 0.859 - KROCC 0.763 0.917 0.883 0.661 - RMSE 1.043 0.256 0.344 11.618 -
@@
TABLE III: Performance comparison of classic and SOTA metrics on different distortion types. The best, second-best, and third-best are highlighted in red, blue, and underlined, respectively.
Method Crit. Distortion types Ds Ns Oc Mx Tc Vc PSNR-p2p PLCC 0.394 0.649 0.525 0.737 0.468 0.208 SROCC 0.333 0.649 0.368 0.673 0.440 0.196 KROCC 0.232 0.482 0.267 0.513 0.311 0.133 RMSE 1.330 0.806 1.082 0.851 1.115 0.893 PSNR-p2pl PLCC 0.296 0.672 0.527 0.666 0.471 0.264 SROCC 0.188 0.668 0.373 0.636 0.437 0.243 KROCC 0.135 0.494 0.270 0.473 0.310 0.166 RMSE 1.383 0.784 1.080 0.939 1.114 0.881 PSNR-YUV PLCC 0.632 0.588 0.561 0.670 0.541 0.393 SROCC 0.617 0.559 0.543 0.659 0.528 0.391 KROCC 0.421 0.404 0.364 0.490 0.369 0.267 RMSE 1.122 0.856 1.052 0.934 1.062 0.840 PSSIM PLCC 0.959 0.771 0.728 0.881 0.464 0.303 SROCC 0.906 0.766 0.564 0.877 0.091 0.166 KROCC 0.719 0.585 0.431 0.691 0.062 0.112 RMSE 0.413 0.675 0.872 0.595 1.118 0.870 PCQM PLCC 0.435 0.900 0.870 0.902 0.718 0.643 SROCC 0.054 0.902 0.872 0.881 0.714 0.623 KROCC 0.047 0.720 0.683 0.691 0.519 0.443 RMSE 1.303 0.462 0.627 0.544 0.879 0.700 MS-PSSIM PLCC 0.703 0.758 0.753 0.880 0.611 0.359 SROCC 0.689 0.737 0.745 0.865 0.565 0.338 KROCC 0.516 0.539 0.551 0.675 0.399 0.233 RMSE 1.029 0.690 0.836 0.597 0.999 0.852 GraphSIM PLCC 0.943 0.885 0.854 0.895 0.628 0.637 SROCC 0.891 0.883 0.844 0.872 0.631 0.619 KROCC 0.685 0.691 0.663 0.670 0.462 0.442 RMSE 0.480 0.492 0.662 0.562 0.983 0.704 MS-GraphSIM PLCC 0.945 0.892 0.877 0.918 0.616 0.673 SROCC 0.893 0.890 0.865 0.898 0.623 0.662 KROCC 0.693 0.701 0.680 0.709 0.459 0.479 RMSE 0.472 0.478 0.610 0.499 0.995 0.675 PHM PLCC 0.950 0.916 0.901 0.941 0.772 0.696 SROCC 0.905 0.909 0.895 0.932 0.769 0.676 KROCC 0.714 0.735 0.716 0.774 0.575 0.488 RMSE 0.452 0.425 0.552 0.425 0.803 0.656 TCDM PLCC 0.943 0.904 0.899 0.952 0.821 0.689 SROCC 0.891 0.903 0.883 0.939 0.818 0.678 KROCC 0.697 0.722 0.710 0.782 0.622 0.493 RMSE 0.481 0.454 0.558 0.385 0.721 0.661 FRSVR (K-fold splicing) PLCC 0.952 0.850 0.911 0.833 0.890 0.733 SROCC 0.902 0.864 0.906 0.829 0.879 0.730 KROCC 0.716 0.678 0.733 0.638 0.654 0.522 RMSE 0.466 0.523 0.511 0.635 0.652 0.621 PointPCA (K-fold splicing) PLCC 0.942 0.847 0.879 0.820 0.842 0.655 SROCC 0.891 0.859 0.885 0.819 0.823 0.671 KROCC 0.703 0.679 0.702 0.634 0.611 0.481 RMSE 0.471 0.517 0.572 0.627 0.711 0.669 MS-ISSM (K-fold splicing) PLCC 0.954 0.876 0.933 0.916 0.902 0.783 SROCC 0.890 0.862 0.923 0.889 0.904 0.754 KROCC 0.695 0.689 0.767 0.699 0.730 0.562 RMSE 0.432 0.512 0.459 0.529 0.545 0.568
V-A Evaluation Criteria
@@
V-B Datasets and PCQA Metrics
To verify the performance of the proposed method across different types of distortions, we use four point cloud subjective datasets for assessment. These datasets include: SJTU [10], WPC [40], M-PCCD[44], and ICIP [45]. Moreover, to optimize the regression model and guarantee strict content separation, we employed a K-fold cross-validation strategy on each dataset[37]. To accommodate the varying sizes of the datasets, the number of folds was explicitly tailored: we utilized 6-fold cross-validation for the ICIP dataset, 8-fold for M-PCCD, 9-fold for SJTU, and 10-fold for WPC. Performance is evaluated using two distinct aggregation methods. The first, denoted as ”K-fold average”, computes the PLCC, SROCC, KROCC, and RMSE metrics per fold and averages them. The second, denoted as ”K-fold splicing”, concatenates the out-of-sample prediction scores for all point clouds across the K-fold iterations to compute the evaluation metrics globally at once [37]. In addition, this paper compares the proposed algorithm with 12 classic and SOTA PCQA metrics. It is worth noting that PointPCA[34] and FRSVR[36] are learning-based methods, and their training and testing follow the same procedures as the proposed method.
@@ V-C Performance Comparison
We evaluate the performance of various PCQA metrics using different datasets. The overall evaluation results for each PCQA metric across these datasets are presented in Table II. To facilitate direct comparison, the best, second-best, and third-best performing metrics are highlighted in red, blue, and underlined text, respectively. Notably, the K-fold average results of the learning-based methods are excluded from the comparison and ranking. Instead, the rankings rely on the global concatenated results using splicing. This approach aggregates predictions from all test folds before calculating the final correlation. It ensures a fair comparison with traditional hand-crafted metrics evaluated on the entire dataset. Specifically, the proposed method ranks highest on the WPC dataset compared to the other PCQA metrics. On the M-PCCD dataset, the proposed method achieves first place in PLCC, SROCC, and KROCC. On the ICIP dataset, it achieves the best performance in PLCC, SROCC, KROCC, and RMSE. Although the proposed method does not achieve the best performance on the SJTU dataset, it closely trails the top-performing TCDM method. On the SJTU dataset, MS-ISSM yields a PLCC of 0.906 and an SROCC of 0.897, compared with the best-performing method, TCDM, which achieves 0.931 and 0.912. Notably, based on the overall rank across all four datasets, the proposed method demonstrates the best performance. It is followed by TCDM, which achieve competitive but lower correlations. Furthermore, compared with other learning-based methods, our method demonstrates better overall performance in both K-fold average and K-fold splicing results. Overall, the proposed method is demonstrated with advanced performance.
Figure 6: Qualitative visual comparisons of distorted point clouds and their objective quality scores. From top to bottom, the models are subjected to octree-based compression (biplanebiplane), trisoup-based compression (sarahsarah), Gaussian noise (flowerpotflowerpot), and video compression (bagbag), with distortion severity increasing from left to right. The scores below each model represent the subjective MOS, predictions from two SOTA methods (TCDM and FRSVR), and our proposed MS-ISSM.
@@
We combine four datasets to validate metric performance across distortion types. Because subjective rating scales differ, we map all scores onto a common scale of 0 to 5 for consistency. Table III presents the results across distortions including octree-based compression (Oc), video-based compression (Vc), trisoup-based compression (Tc), noise (Ns), downsampling (Ds), and mixing (Mix). As with previous tables, the best, second-best, and third-best performing metrics highlighted in red, blue, and underlined text. Overall, the proposed MS-ISSM exhibits strong performance in octree-based compression, trisoup-based compression, and video-based compression. Specifically, it achieves the highest PLCC, SROCC, KROCC, and RMSE in these three categories. For downsampling distortion, MS-ISSM performs competitively, securing the second-best PLCC and RMSE.
@@
Regarding mixing distortion, MS-ISSM performs lower than TCDM. The mixing distortion samples originate entirely from the SJTU dataset, where TCDM parameters were explicitly fitted, providing a distributional advantage. However, in the video-based compression scenario, MS-ISSM maintains a clear lead. While FRSVR obtains an SROCC of 0.730 and an RMSE of 0.621, MS-ISSM reaches a higher SROCC of 0.754 and a lower RMSE of 0.568.
@@
In overall, MS-ISSM achieves the strong performance across different distortion types. This generalization capability confirms the effectiveness of our design in three key aspects. First, the RBF-based implicit representation reconstructs continuous surfaces to resolve sparsity and geometric jitter. This bypasses discrete point-to-point matching errors, yielding superior stability in V-PCC and compression distortions. Second, our hierarchical approach captures both global structural shifts from downsampling and local high-frequency artifacts from compression. Finally, the ResGrouped-MLP combines Log-Modulus transformation for distribution rectification with channel-wise attention for adaptive feature weighting, ensuring robust prediction in mixed distortion scenarios.
@@
Additionally, to further compare the performance of the MS-ISSM method and the RBFIM method in terms of compression distortion, we evaluated datasets containing compression distortion from the three aforementioned datasets. Based on the type of compression distortion, the datasets were divided into G-PCC compression distortion and V-PCC compression distortion. Table IV shows that RBFIM performs better on both subsets of the ICIP dataset. However, MS-ISSM outperforms RBFIM on the G-PCC and V-PCC subsets of the WPC and M-PCCD datasets. Furthermore, MS-ISSM achieves higher correlations and lower errors on the combined ALL dataset. This indicates that the multi-scale implicit feature method provides stable generalization across diverse compression scenarios.
@@
TABLE IV: Comparison of the RBFIM and MS-ISSM (K-fold splicing) methods under different compression distortions.
@@
ICIP-GPCC PLCC 0.993 0.975
SROCC 0.971 0.968
KROCC 0.870 0.846
RMSE 0.020 0.027
ICIP-VPCC PLCC 0.969 0.969
SROCC 0.976 0.956
KROCC 0.895 0.840